# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 4 Adaptive Version
then $M_{\mathcal{F}}^{\operatorname{adloc}}(\bar{f},d/2^{J-3},4(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (4.2) holds for $\varepsilon_{J}$, then we have the following upper bound on $P(A_{J})$:
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{d}{2^{J-3}},4(C+1)\right)\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
This equivalently gives us a lower bound on $P(A_{J}^{c})$ of $1-6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
Now, set $\kappa=\frac{3(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}$. Suppose $J$ satisfies (4.2). Then $J\leq J^{\ast}$ and by definition $J^{\ast}\leq\overline{J}$. Recall also that $f^{\ast}$ was the output after $\overline{J}$ steps. Hence by the Cauchy sequence argument in (2.2), since $J\leq\overline{J}$, we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. If we assume further that $\|\Upsilon_{J}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, then by the triangle inequality, $\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\cdot\varepsilon_{J}$. Now call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, and notice we have $P(A_{J}^{c})\leq P(B_{J})$ for such $J$ satisfying (4.2). We already gave a lower bound on $P(A_{J}^{c})$ for such $J$. Thus, we have proven for all $J$ satisfying (4.2) that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
holds with probability at least
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (4.2) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (4.2) will not hold for all sufficiently large $J$. This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$, $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (4.3)
The first inequality holds trivially for $J=1$ as $\kappa\varepsilon_{1}=3d$ and $d$ is the diameter of $\mathcal{F}$, so both sides of the inequality are 0. Similarly we extend this result to all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$ so the left-hand side of the first inequality is 0.
Next, suppose $x\in[\varepsilon_{J},\varepsilon_{J-1})$. Then
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=2\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
Then integrate:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$

### v5

@@ Appendix C Proofs for Section 4
$\leq 3\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(C,\sigma,L)}},4c)\right]^{2}\cdot\frac{a_{J}}{1-a_{J}}.$
Since $J^{\ast}>1$ and $1\leq J\leq J^{\ast}$, (4.2) implies both that
$\bigg{[}\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(C,\sigma,L)}},4c)\bigg{]}^{2}\leq\exp(n\varepsilon_{J}/2)$
and $\tfrac{1}{1-a_{J}}\leq 2$. Hence
$\mathbb{P}(A_{J})$ $\leq 6\exp(n\varepsilon_{J}^{2}/2)\cdot a_{J}=6\exp(-n\varepsilon_{J}^{2}/2).$
Next, recall we set $f^{\dagger}=\Upsilon_{J^{\dagger}}$ as the output of exactly $J^{\dagger}-1$ steps. Define $B_{j}=\{\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{j}\}$ where $\kappa=\tfrac{1+3c}{2\sqrt{\Lambda(c/2-1,\sigma,L)}}.$
Part 3: Bounding $\mathbb{P}(B_{J})$ for integers $-\infty<J\leq J^{\ast}$.
Well, let us first show $\mathbb{P}(B_{J})\leq\mathbb{P}(A_{J})$ for $1\leq J\leq J^{\ast}$. Suppose $B_{J}$ holds. By comparing (4.1) with (4.2), we see that $J^{\ast}\leq J^{\dagger}$. Hence $J\leq J^{\dagger}$. Thus, the triangle inequality and Lemma 3.3 applied with $\Upsilon_{J}$ and $f^{\dagger}=\Upsilon_{J^{\dagger}}$ imply
$\kappa\varepsilon_{J}$ $<\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}+\|\Upsilon_{J}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{d(2+4c)}{c2^{J}}.$
Then we again conclude $A_{J}$ holds since
$\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{J}-\tfrac{d(2+4c)}{c2^{J}}=\tfrac{d}{2^{J-1}}.$ (C.1)
In other words $B_{J}\subseteq A_{J}$, so
$\mathbb{P}(B_{J})\leq\mathbb{P}(A_{J})\leq 6\exp(-n\varepsilon_{J}^{2}/2),$
since $J\leq J^{\ast}$ and we can also apply the bound from Part 2. Since $\mathbb{P}(B_{J})=0$ for all $J\leq 0$ due to $\kappa\varepsilon_{J}>d$ for such $J$, our bound in (C.1) holds for all integers $-\infty<J\leq J^{\ast}$.
Part 4: Bounding $\mathbb{P}(\|\bar{f}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}x)$ for all $x\geq\varepsilon_{J^{\ast}}$, where $\kappa^{\prime}=(2+\tfrac{1+2c}{1+3c})\kappa$.
We now re-use the logic from Part 4 of the proof of the non-adaptive Theorem 3.13 but using $f^{\dagger}$ and $f^{\ddagger}$ in place of $f^{\ast}$ and $f^{\ast\ast}$, respectively. Recall $f^{\ddagger}=\Upsilon_{J^{\ddagger}+1}$ is the output of $J^{\ddagger}\geq J^{\dagger}$ steps. Pick any $x\geq\varepsilon_{J^{\ast}}$ so for some $0<J\leq J^{\ast}$ we have $x\in[\varepsilon_{J},\varepsilon_{J-1})$. Then we have
$\|\bar{f}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\|f^{\dagger}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}$
$=\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\|\Upsilon_{J^{\ddagger}+1}-\Upsilon_{J^{\dagger}}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{d(2+4c)}{c2^{J^{\dagger}}}$
$=\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa\varepsilon_{J^{\dagger}}$
$\leq\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa x.$
The last line used that $x\geq\varepsilon_{J^{\ast}}\geq\varepsilon_{J^{\dagger}}$, which in turn follows from $J^{\ast}\leq J^{\dagger}$. Using the definition $\kappa^{\prime}$ and that $2\kappa x\geq\kappa\varepsilon_{J-1}$ since $x\geq\varepsilon_{J}=\varepsilon_{J-1}/2$, we have
$\mathbb{P}(\|\bar{f}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}x)$ $\leq\mathbb{P}(\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa x>\kappa^{\prime}x)$
$=\mathbb{P}(\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}>2\kappa x)$
$\leq\mathbb{P}(\|\bar{f}-f^{\dagger}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{J-1})$
$=\mathbb{P}(B_{J-1})$
$\leq 6\exp(-n\varepsilon_{J-1}^{2}/2)$
$\leq 6\exp(-nx^{2}/2).$
The last few lines used Part 3’s bound on $\mathbb{P}(B_{J})$ and that $X<\varepsilon_{J-1}$.
Part 5: Bounding $\mathbb{E}_{\vec{X},\vec{Y}}\|\bar{f}-f^{\ddagger}(\vec{X},\vec{Y})\|_{L_{2}(\mathbb{P}_{X})}^{2}$.
We integrate Part 4’s bound:
$\mathbb{E}_{\vec{X},\vec{Y}}\|\bar{f}-f^{\ddagger}(\vec{X},\vec{Y})\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$=\int_{0}^{\infty}\mathbb{P}(\|\bar{f}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}^{2}>x)\mathrm{d}x$
$=2{\kappa^{\prime}}^{2}\int_{0}^{\infty}t\cdot\mathbb{P}(\|\bar{f}-f^{\ddagger}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}t)\mathrm{d}t$